OpenStudy (anonymous):

log (base 3)27.. solve the log?

OpenStudy (asadkarim7):

3

OpenStudy (anonymous):

3^x=27 is the same problem so 3

OpenStudy (anonymous):

27 = 3^x so x = 3

OpenStudy (anonymous):

Thanks guys! im on a timed test so any help would be greatly appreciated!! :D

OpenStudy (anonymous):

no probs